OpenStudy (anonymous):

Given a = -3, b = 4 and c = -5, evaluate a - b - c. -12 -6 -2

OpenStudy (anonymous):

a-b-c= -3-4-(-5)= -3-4+5=-7+5=-2 Answer

OpenStudy (anonymous):

thank you thank you thank you!:)